OpenStudy (anonymous):

i need some help i have to create an equivalent system of equations with 2x + 7y = 1 -3x – 4y = 5

OpenStudy (anonymous):

can someone help me i have been stuck on this for almost 2 hours

OpenStudy (anonymous):

6x+21y=3 -6x-8y=10 13y=13 y=1

OpenStudy (anonymous):

There are a lot of equivalent systems. This is a strange question to ask, but I suppose the best approach to take would be to solve the equations first. Shadowfenix has done the bulk of that for you :p

OpenStudy (anonymous):

Well i dont really understand how to work through the problem.

OpenStudy (anonymous):

Which part are you stuck with?